Two equal positive charges q1 = q2 = 4.0 μC are located at x = 0, y = 0.30 m, x = 0, y = -0.30 m and respectively. What are the magnitude and direction of the total electric force that q1 and q2 exert on a third charge at Q = 4.0 μC x = 0.40 m, y = 0?


1
Expert's answer
2022-02-16T08:35:45-0500

Answer

Force on third charge Q is only horizontal components addition and vertical cancelled out.

"F=2*\frac{kq_1q_2}{r^2}\\=\frac{2*16*10^{-12}}{0.5}\\=64**10^{-12}N"